According to a survey only 20% of customers who visited tje website of a major retail store made a purchase. random sample sizes of 75 are selected.
What proportion of the sample will have between 10% and 40% of customers who will make a purchase after visiting the website

Answer by Boreal(15235)

the mean is np=75*0.20=15
the variance is np(1-p)=15*0.8=12
the sd is sqrt(V)=3.464
10% is 7.5 customers (we would consider 8 the bottom number)
40% is 30 customers
with the continuity correction factor, we would use 7.5 and 30.5 for the two values.
z=(7.5-15)/3.464 with z of -2.165
for 30.5
z=(30.5-15)/3.464
=+4.47
the probability of z's being in between those values is 98.5%